The sides of a square are lengthened by 7 cm, the area becomes 256cm^2. Find the length of the original square.
deb_adkins2@yahoo.com

Answer by ankor@dixie-net.com(22740)

The sides of a square are lengthened by 7 cm, the area becomes 256cm^2.
Find the length of the original square.
:
Let x = the length of the original square
then
(x+7) = the length of the larger square
The area
(x+7)^2 = 256
FOIL
x^2 + 14x + 49 = 256
:
x^2 + 14x + 49 - 256 = 0
:
x^2 + 14x - 207 = 0
Factors to
(x+23)(x-9) = 0
Positive solution is all we want here
x = 9 cm the side of the original square
:
:
Check 16^2 = 256